# Stochastic Conditional Diffusion Models for Robust Semantic Image Synthesis

### v2

@@ Abstract
Semantic image synthesis (SIS) is a task to generate realistic images corresponding to semantic maps (labels). It can be applied to diverse real-world practices such as photo editing or content creation. However, in real-world applications, SIS often encounters noisy user inputs. To address this, we propose Stochastic Conditional Diffusion Model (SCDM), which is a robust conditional diffusion model that features novel forward and generation processes tailored for SIS with noisy labels. It enhances robustness by stochastically perturbing the semantic label maps through Label Diffusion, which diffuses the labels with discrete diffusion. Through the diffusion of labels, the noisy and clean semantic maps become similar as the timestep increases, eventually becoming identical at $t=T$. This facilitates the generation of an image close to a clean image, enabling robust generation. Furthermore, we propose a class-wise noise schedule to differentially diffuse the labels depending on the class. We demonstrate that the proposed method generates high-quality samples through extensive experiments and analyses on benchmark datasets, including a novel experimental setup simulating human errors during real-world applications.
@@ 1 Introduction
  • We propose Stochastic Conditional Diffusion Model, a novel and robust conditional diffusion model for SIS with a discrete forward process of labels and a continuous reverse process of images.
  • We design Label Diffusion, a discrete diffusion process for labels, which enables differential conditioning on the semantic labels.
@@ 2 Related Works
Since Pix2pix (Isola et al., 2017) have established a general framework for SIS, conditional Generative Adversarial Networks (GANs) are widely used in SIS (Yang et al., 2019; Zhu et al., 2017; Ntavelis et al., 2020; Tang et al., 2020b; Tan et al., 2021a, b; Shi et al., 2022). SPADE (Park et al., 2019) proposes spatially-adaptive normalization and successfully preserves semantic information. RESAIL (Shi et al., 2022) proposes retrieval-based spatially adaptive normalization, and OASIS (Sushko et al., 2020) designs the discriminator as a semantic segmentation network. SAFM (Lv et al., 2022) proposes shape-aware position descriptors to modulate the features. Recently, diffusion models (DMs) also have been proposed for SIS. SDM (Wang et al., 2022b) encodes the semantic label map with SPADE. LDM (Rombach et al., 2022) leverages a latent space for the conditions including the semantic maps. PITI (Wang et al., 2022a) pre-trains the semantic latent space and finetunes it with the RGB-preprocessed semantic mask images, rather than using the maps with class indexes like most studies including ours. FLIS (Xue et al., 2023) incorporates not only semantic label maps but also additional text inputs. Most of these works have recognized the applicability of SIS in real-world scenarios. However, to the best of our knowledge, our method is the first DM-based model to address the issue of noisy user inputs in SIS.
@@
By modifying the U-Net (Ronneberger et al., 2015) architecture to incorporate the conditions into the network, conditional diffusion models are prevalently leveraged for conditional generations. ADM (Dhariwal & Nichol, 2021), for example, encodes the class embedding into the network with AdaGN and utilizes classifier guidance. LDM (Rombach et al., 2022) extracts features from the various conditions and encodes them into the network through concatenation or cross-attention. Meanwhile, SDM (Wang et al., 2022b) replaces GroupNorm of U-Net decoder with SPADE to embed the labels into the network in a spatially adaptive manner. UNIT-DDPM (Sasaki et al., 2021) uses two different diffusion models and a domain translation function for the training and sampling of an image-to-image translation model. Our method diffuses labels with a carefully designed discrete diffusion process and generates images conditioned on the diffused labels, formulating a novel conditional diffusion model.
@@ 4 Stochastic Conditional Diffusion Model
Figure 2: Generation process of SCDM. The Stochastic Conditional Diffusion Model (SCDM) is a robust conditional diffusion model for semantic image synthesis. SCDM consists of a discrete forward process for labels and a continuous reverse process for images. It improves the robustness to noisy semantic labels as well as generation performance on clean semantic labels. $\boldsymbol{z}_{0}$ denotes the $i$-th pixel of the semantic map, i.e., $\boldsymbol{z}_{0}=\mathbf{y}_{0}^{i}$ where $\mathbf{y}_{0}=\{\mathbf{y}_{0}^{1},...,\mathbf{y}_{0}^{H\times W}\}$
@@ 4.2.1 Transition Matrix for Label Masking
To gradually erase the information of the semantic map and increase similarity among different maps as $t=0\rightarrow t=T$, we designed our Label Diffusion process by progressively masking labels. In other words, the original semantic labels are converted into the absorbing state ‘masked’ with some probability at each timestep. Consequently, all semantic maps eventually become identical at $t=T$, each filled with ‘masked’ at every pixel.
Given $C$ semantic classes and one masked state, we define the transition matrix $\boldsymbol{Q}_{t}\in\mathbb{R}^{(C+1)\times(C+1)}$ at a timestep $t$ as
@@
where the absorbing state is added as $(C+1)$-th class and $\beta_{t,c}$ is the probability of a label of class $c$ to be masked. Note that class-wise defined probability $\beta_{t,c}$ enables class-wise noise scheduling, which will be further discussed in the following section. We assume that $\boldsymbol{Q}_{t}$ is applied to each pixel of the semantic map independently. Since the ‘masked’ state is the absorbing state, once the state is in $C+1$ then the next state is always $C+1$.
@@ 4.2.2 Noise Scheduling
We observe that the semantic information of small objects in an image is prone to be lost at a relatively early stage of diffusion compared to larger objects. Moreover, for a rare class of objects in the dataset, it would be hard to learn their semantics if their labels are masked quickly. Thus, we propose a class-wise noise schedule to differentially transform semantic labels depending on the class. We designed $\gamma_{t,c}$ to ensure that labels occupying smaller areas and rarely appearing in the dataset are transitioned into the absorbing state more slowly and at a later time. The class-wise schedule improves generation quality for small and rare objects. For more details, see Section 6.2 and Appendix F.4.
@@ 4.3 Training and Sampling
where $\alpha_{t}$ determines the noise level for input image $\mathbf{x}$ at timestep $t$, $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a Gaussian noise, and $\lambda$ is a balancing hyperparameter. This is similar to the hybrid loss (Nichol & Dhariwal, 2021), with a slight adaptation of using $\mathbf{y}_{t}$. See Appendix A.1 for a detailed derivation of the objective function.
To generate a sample, by the definition of the generation process, we first sample $\mathbf{y}_{1:T}$ from the semantic map $\mathbf{y}_{0}$ and feed it sequentially to the process. In addition, we formulate the classifier-free guidance (Ho & Salimans, 2021) in our model:
@@
where the constant extrapolation scale is denoted by $w$. Following extrapolation, we apply dynamic thresholding (Saharia et al., 2022) and subsequently randomly sample $\mathbf{x}_{t-1}$ utilizing $\mathbf{x}_{t}$ and $\tilde{\mathbf{x}}_{0}^{(t)}$. The complete training and sampling algorithm is in Appendix B.
@@ 4.4 Discussion
[…]
@@
The proof is available in Appendix F.5. Proposition 2 implies that the expectation of implicit classifier gradients in our method over all possible $\mathbf{y}_{t}$ given $\mathbf{y}_{0}$ is equivalent to $\nabla_{\mathbf{x}_{t}}\log q(\mathbf{y}_{0}|\mathbf{x}_{t})$ after time-dependent scaling. First, the scaling factor $(1-\gamma_{t})$ starts from 0 when $t=T$ and is set to 1 when $t=0$. In other words, our method acts like an unconditional generation at the beginning of the reverse process and gets stronger guidance as $t$ goes to 0. Second, the expectation of the classifier gradient in our method is the same direction as the one in the baseline with fixed class labels. Note that this does not mean that our method has the same guidance as the baseline with fixed labels and time-dependent scaling. For more discussion, see Appendix F.5.
@@ 5.1 Noisy SIS benchmark
We evaluate our method on ADE20K (Zhou et al., 2017) dataset. ADE20K contains 20K images for training and 2K images for test annotated with 151 classes including the ‘unlabeled’ class. Additionally, we introduce three new experimental setups to assess generation performance under noisy conditions using the ADE20K dataset as follows:
@@
[Edge] This setup masks the edges of instances with an unlabeled class. This setup imitates incomplete annotations around edges, especially between instances. We observed that human annotators occasionally leave the pixels on the boundaries of instances as ‘unlabeled’ due to their inherent ambiguity, see Appendix E for examples. Assuming a pixel with a different class compared to its neighbor is the edge of the instance, we detect edges using the label map. Then, we fill the semantic map pixels with a distance of 2 or less from the edges with the unlabeled class.
@@
More experimental results with other benchmark datasets (e.g., CelebAMask-HQ (Lee et al., 2020) and COCO-Stuff (Caesar et al., 2018)) are in Appendix F.
@@ 5.3 SIS with Noisy Labels
Table 1 shows the performances of our method and baselines under the three settings. SCDM demonstrates its superior robustness to all three types of noise in generation quality measured by FID, compared to other baselines. Notably, in the three setups, the performance gaps between the best baseline scores and our method are +2.1, +6.4, and +8.2, respectively. We also evaluate the semantic correspondence between the clean ground-truth label maps and the generation results and report mIoU scores. Our method achieved the best mIoU performance among DM-based models in all three settings. Compared to strong GAN-based baselines, including SAFM (Lv et al., 2022) that leverages extra ground truth instance maps during sampling, our results show comparable results in correspondence.
@@ 6.1 Ablation Study
Figure 4: Generation results with and without Label Diffusion. The results are sampled with the fixed random seeds and same $\mathbf{x}_{T}$, and generated with (a) clean labels, (b) DS, (c) Edge, and (d) Random setup noisy labels, respectively.
@@ 6.2 Effect of Class-wise Noise Schedule
In this section, we elucidate the effect of our class-wise noise schedule by comparing the generation results of two different models using the class-wise schedule and linear and uniform schedule, i.e., $\eta\rightarrow 0$. We observe that the class-wise schedule clearly improves the image quality, especially in terms of semantic correspondence (mIoU ($\uparrow$) of 49.4 (class-wise) $>$ 43.0 (uniform) on ADE20K). The quantitative and qualitative results are presented in the Appendix F.4. Furthermore, the class-wise schedule exhibits superiority in small and rare class synthesis. By organizing the classes into three groups - frequent, common, and rare - based on their $\psi_{c}\phi_{c}$ scores, we compare mIoU per each group on ADE20K and report the performance in Table 4. Notably, our class-wise schedule exhibited mIoU gain in all the groups, with the highest performance improvement of +8.1 in the rare group.

### v3

@@ Abstract
Semantic image synthesis (SIS) is a task to generate realistic images corresponding to semantic maps (labels). However, in real-world applications, SIS often encounters noisy user inputs. To address this, we propose Stochastic Conditional Diffusion Model (SCDM), which is a robust conditional diffusion model that features novel forward and generation processes tailored for SIS with noisy labels. It enhances robustness by stochastically perturbing the semantic label maps through Label Diffusion, which diffuses the labels with discrete diffusion. Through the diffusion of labels, the noisy and clean semantic maps become similar as the timestep increases, eventually becoming identical at $t=T$. This facilitates the generation of an image close to a clean image, enabling robust generation. Furthermore, we propose a class-wise noise schedule to differentially diffuse the labels depending on the class. We demonstrate that the proposed method generates high-quality samples through extensive experiments and analyses on benchmark datasets, including a novel experimental setup simulating human errors during real-world applications. Code is available at https://github.com/mlvlab/SCDM.
@@ 1 Introduction
  • We propose Stochastic Conditional Diffusion Model (SCDM), a robust conditional diffusion model for SIS that incorporates Label Diffusion, a discrete diffusion process for labels that enables differential conditioning on semantic labels.
@@ 2 Related Works
Since Pix2pix (Isola et al., 2017) have established a general framework for SIS, conditional Generative Adversarial Networks (GANs) are widely used in SIS (Yang et al., 2019; Zhu et al., 2017; Ntavelis et al., 2020; Tang et al., 2020b; Tan et al., 2021a, b; Shi et al., 2022). SPADE (Park et al., 2019) proposes spatially-adaptive normalization and successfully preserves semantic information. Since SPADE, many normalization-based approaches have been presented (Tan et al., 2021a; Lv et al., 2022). For instance, CLADE (Tan et al., 2021b) adopts class-adaptive normalization and RESAIL (Shi et al., 2022) proposes retrieval-based spatially adaptive normalization, and OASIS (Sushko et al., 2020) designs the discriminator as a semantic segmentation network. INADE (Tan et al., 2021a) utilizes class-level conditional modulation, and SAFM (Lv et al., 2022) proposes shape-aware position descriptors to modulate the features.
Recently, diffusion models (DMs) also have been proposed for SIS. SDM (Wang et al., 2022b) encodes the semantic label map with SPADE. LDM (Rombach et al., 2022) leverages a latent space for the conditions including the semantic maps. PITI (Wang et al., 2022a) pre-trains the semantic latent space and finetunes it with the RGB-preprocessed semantic mask images, rather than using the maps with class indexes like most studies including ours. FLIS (Xue et al., 2023) incorporates not only semantic label maps but also additional text inputs. Most of these works have recognized the applicability of SIS in real-world scenarios. However, to the best of our knowledge, our method is the first DM-based model to address the issue of noisy user inputs in SIS.
@@
By modifying the U-Net (Ronneberger et al., 2015) architecture to incorporate the conditions into the network, conditional diffusion models are prevalently leveraged for conditional generations. ADM (Dhariwal & Nichol, 2021), for example, encodes the class embedding into the network with AdaGN and utilizes classifier guidance. LDM (Rombach et al., 2022) extracts features from the various conditions and encodes them into the network through concatenation or cross-attention. DDMI (Park et al., 2024) also generates data with latent diffusion but in a domain-agnostic manner, by linking data to a continuous function within a shared latent space. Meanwhile, SDM (Wang et al., 2022b) replaces GroupNorm of U-Net decoder with SPADE to embed the labels into the network in a spatially adaptive manner. UNIT-DDPM (Sasaki et al., 2021) uses two different diffusion models and a domain translation function for the training and sampling of an image-to-image translation model. Recently, T2I diffusion models such as FLIS (Xue et al., 2023), ControlNet (Zhang et al., 2023), and GLIGEN (Li et al., 2023) have paved the way for adding spatial controls to large pretrained diffusion models. Our method diffuses labels with a carefully designed discrete diffusion process and generates images conditioned on the diffused labels, formulating a novel conditional diffusion model.
@@ 4 Stochastic Conditional Diffusion Model
Figure 2: Generation process of SCDM. The Stochastic Conditional Diffusion Model (SCDM) is a robust conditional diffusion model for semantic image synthesis. SCDM consists of a discrete forward process for labels and a continuous reverse process for images. It improves the robustness to noisy semantic labels as well as generation performance on clean semantic labels. $\boldsymbol{z}_{0}$ denotes the $i$-th pixel of the semantic map, i.e., $\boldsymbol{z}_{0}=\mathbf{y}_{0}^{i}$ where $\mathbf{y}_{0}=\{\mathbf{y}_{0}^{1},...,\mathbf{y}_{0}^{H\times W}\}$.
@@ 4.1 Definitions
For simplicity, we employed the same $t$ for both the image $\mathbf{x}$ and the label $\mathbf{y}$, but the noise levels are separately controlled. We manually designed a different noise schedule for the labels, referred to as the class-wise noise schedule, which will be discussed in the following section. Consequently, the diffusion processes for $\mathbf{x}$ and $\mathbf{y}$ are not synchronized regarding noise levels, and the synchronization between them is not necessary.
@@ 4.2.1 Transition Matrix for Label Masking
To gradually erase the information of the semantic map and increase similarity among different maps as $t=0\rightarrow t=T$, we designed our Label Diffusion process by progressively masking labels. In other words, the original semantic labels are converted into the absorbing state (i.e., [mask]) with some probability at each timestep. Consequently, all semantic maps eventually become identical at $t=T$, each filled with [mask] at every pixel. In addition, as the semantic labels are discrete variables, i.e., classes, it is natural to adopt the discrete diffusion.
Given semantic classes $C$ and absorbing state [mask], we define the transition matrix $\boldsymbol{Q}_{t}\in\mathbb{R}^{(C+1)\times(C+1)}$ at a timestep $t$ as:
@@
where the absorbing state is added as $(C+1)$-th class and $\beta_{t,c}$ is the probability of a label of class $c$ to be masked. Note that class-wise defined probability $\beta_{t,c}$ enables class-wise noise scheduling, which will be further discussed in the following section. We assume that $\boldsymbol{Q}_{t}$ is applied to each pixel of the semantic map independently. Since [mask] is the absorbing state, once the state is in $C+1$ then the next state is always $C+1$.
@@ 4.2.2 Noise Scheduling
We observe that the semantic information of small objects in an image is prone to be lost at a relatively early stage of diffusion compared to larger objects. Moreover, for a rare class of objects in the dataset, it would be hard to learn their semantics if their labels are masked quickly. Thus, we propose a class-wise noise schedule to differentially transform semantic labels depending on the class. We designed $\gamma_{t,c}$ to ensure that labels occupying smaller areas and rarely appearing in the dataset are transitioned into the absorbing state more slowly and at a later time. The class-wise schedule improves generation quality for small and rare objects. For more details, see Section 6.2 and Appendix G.4.
@@ 4.3 Training and Sampling
where $\alpha_{t}$ determines the noise level for input image $\mathbf{x}$ at timestep $t$, $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a Gaussian noise, and $\lambda$ is a balancing hyperparameter. This is similar to the hybrid loss (Nichol & Dhariwal, 2021), with a slight adaptation of using $\mathbf{y}_{t}$. See Appendix A.1 for a detailed derivation of the objective function. Note that Label Diffusion does not significantly impact the training cost of the main diffusion.
To generate a sample, by the definition of the generation process, we first sample $\mathbf{y}_{1:T}$ from the semantic map $\mathbf{y}_{0}$ and feed them sequentially to the process. In addition, we formulate the classifier-free guidance (Ho & Salimans, 2021) in our model:
@@
where the constant extrapolation scale is denoted by $w$. Following extrapolation, we apply dynamic thresholding (Saharia et al., 2022) and subsequently randomly sample $\mathbf{x}_{t-1}$ utilizing $\mathbf{x}_{t}$ and $\tilde{\mathbf{x}}_{0}^{(t)}$.
The complete training and sampling algorithm is in Appendix B. Note that the overall algorithms remain mostly unchanged and the only modification involves incorporating [mask] into labels through Label Diffusion. This introduces minimal computational overhead.
@@ 4.4 Discussion
Suppose there exists a differentiable function $f_{i}$ such that $q(\mathbf{y}_{0}^{i}|\mathbf{x}_{t})=\text{Cat}(\mathbf{y}_{0}^{i};p=f_{i}(\mathbf{x}_{t}))$ and $\mathbf{y}_{0}^{1}|\mathbf{x}_{t},...,\mathbf{y}_{0}^{H\times W}|\mathbf{x}_{t}$ are independent, where $i\in\{1,...,H\times W\}$ denotes the index of a pixel in semantic map and $\mathbf{y}_{0}^{i}\in\mathbb{R}^{C+1}$ is a one-hot vector. With $q(\mathbf{y}_{t}^{i}|\mathbf{y}_{0}^{i})$ from Eq. (13) and $\gamma_{t,c}=\gamma_{t}\text{ for any }c$, we have the following relationship;
@@
The proof is available in Appendix G.5. Proposition 2 implies that the expectation of implicit classifier gradients in our method over all possible $\mathbf{y}_{t}$ given $\mathbf{y}_{0}$ is equivalent to $\nabla_{\mathbf{x}_{t}}\log q(\mathbf{y}_{0}|\mathbf{x}_{t})$ after time-dependent scaling. First, the scaling factor $(1-\gamma_{t})$ starts from 0 when $t=T$ and is set to 1 when $t=0$. In other words, our method acts like an unconditional generation at the beginning of the reverse process and gets stronger guidance as $t$ goes to 0. Second, the expectation of the classifier gradient in our method is the same direction as the one in the baseline with fixed class labels. Note that this does not mean that our method has the same guidance as the baseline with fixed labels and time-dependent scaling. For more discussion, see Appendix G.5.
@@ 5.1 Noisy SIS benchmark
We evaluate our method based on ADE20K (Zhou et al., 2017) dataset. ADE20K contains 20K images for training and 2K images for test annotated with 151 classes including the ‘unlabeled’ class. Additionally, we introduce three new experimental setups to assess generation performance under noisy conditions using the ADE20K dataset as follows:
@@
[Edge] This setup masks the edges of instances with an unlabeled class. This setup imitates incomplete annotations around edges, especially between instances. We observed that human annotators occasionally leave the pixels on the boundaries of instances as ‘unlabeled’ due to their inherent ambiguity, see Appendix F for examples. Assuming a pixel with a different class compared to its neighbor is the edge of the instance, we detect edges using the label map. Then, we fill the semantic map pixels with a distance of 2 or less from the edges with the unlabeled class.
@@
More experimental results with other benchmark datasets (e.g., CelebAMask-HQ (Lee et al., 2020) and COCO-Stuff (Caesar et al., 2018)) are in Appendix G.
@@ 5.3 SIS with Noisy Labels
Table 1 shows the performances of our method and baselines under the three settings. SCDM demonstrates its superior robustness to all three types of noise in generation quality measured by FID, compared to other baselines. Notably, in the three setups, the performance gaps between the best baseline scores and our method are +2.1, +6.4, and +8.2, respectively. We also evaluate the semantic correspondence between the clean ground-truth label maps and the generation results and report mIoU scores. Our method achieved the best mIoU performance among DM-based models in all three settings. Compared to strong GAN-based baselines, including SAFM (Lv et al., 2022) (denoted with ‘*’ in Table 1) that leverages extra ground truth instance maps during sampling, our results show comparable results in correspondence.
@@
5.4 SIS with Clean Labels (Standard SIS)
Table 2: Quantitative performance comparison on generation quality. The baseline methods are categorized into Generative Adversarial Networks (GAN) and Diffusion Models (DM). For FID, lower is better. For LPIPS and mIoU, higher is better. The best results among the diffusion-based approaches are boldfaced, and the best results overall are underlined. ‘-’ indicates that the method did not report the metric or train the dataset, or the checkpoint or samples are not publicly available. ‘Seg’ denotes that the method leverages a pre-trained segmentation network during training. ‘†’ denotes a one-shot method.
Methods Seg CelebAMask-HQ ADE20K COCO-Stuff
GAN RESAIL (Shi et al., 2022) - - - 30.2 - 49.3* 18.3 - 44.7
SAFM (Lv et al., 2022) - - - 32.8 - 52.6 24.6 - 43.3
ECGAN (Tang et al., 2023a) - - - 25.8 0.52 50.6* 15.7 - 46.3
ECGAN++ (Tang et al., 2023b) - - - 24.7 0.54 52.7* 14.9 - 47.9
pix2pixHD (Wang et al., 2018) 38.5 0 76.1 81.8 0 20.3* 111.5 0 14.6
SPADE (Park et al., 2019) 29.2 0 75.2 33.9 0 44.5 33.9 0 36.9
CC-FPSE (Liu et al., 2019) - - - 31.7 0.078 47.3 19.2 0.098 40.8
DAGAN (Tang et al., 2020a) 29.1 0 76.6 31.9 0 45.5 - - -
GroupDNet (Zhu et al., 2020b) 25.9 0.365 76.1 41.7 0.230 33.7 - - -
OASIS (Sushko et al., 2020) - - - 28.3 0.286 50.9 17.0 0.328 44.2
INADE (Tan et al., 2021a) 21.5 0.415 74.1 35.2 0.459 41.4 - - -
SCGAN (Wang et al., 2021) 20.8 0 75.5 29.3 0 50.0 18.1 0 41.7
CLADE (Tan et al., 2021b) 30.6 0 75.4 35.4 0 44.7 29.2 0 36.9
Unconditional Generator† (Chae et al., 2024) 18.5 - 53.1 - - - - - -
LDM (Rombach et al., 2022) 21.5 0.315 74.6 36.5 0.417 23.2 - - -
PITI (Wang et al., 2022a) - - - 27.3 - - 15.8 0.489 32.2
Ours 17.4 0.418 77.2 26.9 0.530 49.4 15.3 0.519 38.1
We also evaluate our method in a standard SIS setting with clean labels. In this experiment, we additionally adopt LPIPS (Zhang et al., 2018) as a diversity metric. Table 2 shows that SCDM achieves the best performance in all three metrics (e.g., FID, LPIPS, mIoU) compared to recent DM-based baselines in all datasets. Also, including GAN-based models, the proposed method shows comparable performances. Specifically, our method surpasses all baselines on CelebAMask-HQ in all three metrics with a significant gain of +1.1 (FID) compared to the state-of-the-art method. Qualitative comparisons and more detailed analysis are in Appendix G.1.
@@ 6.1 Ablation Study
Figure 4: Generation results with and without Label Diffusion. The results are sampled with the fixed random seeds and the same $\mathbf{x}_{T}$, and generated with (a) clean labels, (b) DS, (c) Edge, and (d) Random setup noisy labels, respectively.
@@ 6.2 Effect of Class-wise Noise Schedule
In this section, we elucidate the effect of our class-wise noise schedule by comparing the generation results of two different models using the class-wise schedule and linear and uniform schedule, i.e., $\eta\rightarrow 0$. We observe that the class-wise schedule clearly improves the image quality, especially in terms of semantic correspondence (mIoU ($\uparrow$) of 49.4 (class-wise) $>$ 43.0 (uniform) on ADE20K). The quantitative and qualitative results are presented in the Appendix G.4. Furthermore, the class-wise schedule exhibits superiority in small and rare class synthesis. By organizing the classes into three groups - frequent, common, and rare - based on their $\psi_{c}\phi_{c}$ scores, we compare mIoU per each group on ADE20K and report the performance in Table 5. Notably, our class-wise schedule exhibited mIoU gain in all the groups, with the highest performance improvement of +8.1 in the rare group.